# RF4D:Neural Radar Fields for Novel View Synthesis in Outdoor Dynamic Scenes

### v2

@@ 6 Conclusion
In this paper, we present RF4D, the first radar-based neural field framework specifically designed for novel view synthesis in outdoor dynamic scenes. To capture the inherent motion in such environments, RF4D incorporates time as an explicit input and predicts two physically grounded quantities, occupancy and radar cross section (RCS), for each query position. These predictions are then used within a radar-specific power rendering formulation, which respects radar sensing physics and enables realistic and physically interpretable measurement synthesis. Extensive experiments across dynamic driving scenarios demonstrate that RF4D consistently outperforms existing methods in both measurement synthesis and occupancy estimation accuracy.
Limitation.
While RF4D is physically grounded and achieves strong performance in synthesizing realistic radar measurements and estimating scene occupancy, its training process remains resource-intensive. Each scene is trained independently, typically requiring around five minutes, which limits its applicability in real-time scenarios. Moreover, handling longer sequences or larger-scale environments significantly increases memory and computation requirements, presenting challenges for deployment in resource-constrained settings.

### v3

@@ 6 Conclusion
In this work, we introduced RF4D, a radar-based neural field framework for novel view synthesis in dynamic outdoor scenes. By integrating temporal modeling with a dedicated scene flow module, RF4D enforces temporal coherence and accurately captures moving targets. The proposed radar-specific power rendering further aligns the neural field representation with radar sensing physics, enhancing both accuracy and interpretability. Comprehensive experiments on public datasets demonstrate that RF4D substantially outperforms existing methods in radar measurement synthesis and occupancy estimation, particularly under dynamic and adverse conditions. These results highlight the potential of radar-based neural fields for robust 4D scene understanding.
Acknowledgment
This work was carried out at the Rapid-Rich Object Search (ROSE) Lab, School of Electrical & Electronic Engineering, Nanyang Technological University (NTU), Singapore.